Find the minimum and maximum wavelengths of sound in water that is in the audible range (20-20000 Hz) for an average human ear. Speed of sound in water = 1450 m s-1.

Grade:upto college level

3 Answers

Aditi Chauhan
9 years ago
Sol. a) for minimum wavelength n = 20 KHz ⇒v = nλ⇒ λ = (1450//20 * 10^3) = 7.25 cm. b) for maximum wavelength n should be minimum ⇒ v = nλ⇒ λ = v/n ⇒ 1450 / 20 = 72.5 m.
Ammar Shaheen
6 years ago
The range of human hearing is20Hz to 20KHz.if lowest audible wavelength is 16.6 find the maximum audible wavelength
Devasi Bheemaram
6 years ago
Maximum wavelength = velocity of sound/maximum audible sound =1450/20 = 72.5 m Minimum wavelength = velocity of sound/minimum audible sound=1450/20000=0.725 m